# Adaptive path planning for efficient object search by UAVs in agricultural fields

### v1

@@ 2.2.1 Monte-Carlo dropout
Dropout is a regularization technique in neural networks that randomly drops connections between neurons, and is regularly applied to avoid over-fitting during training by forcing the network to learn different ’routes’ through the network to get to the same outcome (Srivastava et al., 2014). Gal and Ghahramani (2016) introduced Monte-Carlo dropout during inference and show that it can be used to assess model uncertainty. Each connection where dropout is applied has the probability of $q$ to be removed, and this process is repeated for $N$ runs in the network. When the network produces the same output during all runs while randomly removing some connections, it indicates that the network is certain. However, when every run results in a different output, the model is uncertain how to handle the input.
Figure 4 shows the architecture of YOLOv8, indicating the backbone, the neck and head of the network and the several connections that pass-through information from earlier layers in the network. The locations where dropout is applied are indicated with crosses. The dropout is applied between the neck and the head of the network, which enables us to reuse the output of the backbone and neck during inference and save computational time.
Figure 4: Schematic (simplified) overview of the YOLOv8 architecture with the locations indicated where Monte-Carlo dropout is applied between the neck and head of the network. Note that the ’conv’, ’C2F’, ’SPPF’ and ’detect’ blocks consists of different combinations of convolution layers, normalization layers, upsampling layers and activation functions. See Jocher et al. (2023) for more details.
@@
where $D_{j,i}$, is the detection of object $i$ in the MC run $j$. $D_{j,i}$ contains the bounding box coordinates, class probabilities and confidence score. To combine the MC detections with a the reference detection $D_{i}^{\textrm{ref}}$ for each forward pass $j$, we calculate the intersection-over-union with all MC detections $D_{j,i}^{\textrm{mc}}$:
$IoU(D_{i}^{\textrm{ref}},D_{j,i}^{\textrm{mc}})=\frac{|D_{i}^{\textrm{ref}}\cap D_{j,i}^{\textrm{mc}}|}{|D_{i}^{\textrm{ref}}\cup D_{j,i}^{\textrm{mc}}|}.$ (6)
When the IoU is larger than 0.5, $D_{j,i}^{\textrm{mc}}$ is added to the detection set $S_{i}$. Note that combining the reference detection with the MC detections is only based on their bounding box coordinates and is class independent.
@@ 2.2.2 Certainty calculation
The average YOLO certainty, $c_{i}^{\textrm{yolo}}$ of reference detection $i$ is calculated by taking the mean of the YOLO confidence values corresponding the detection set $S_{i}$:
@@
$c_{i}^{\textrm{loc}}=\frac{1}{\|S_{i}\|}\cdot\sum_{D_{j,i}^{\textrm{mc}}\in S_{i}}IoU(D_{i}^{\textrm{ref}},D_{j,i}^{\textrm{mc}}).$ (9)
The localization certainty is based on work of Blok et al. (2022). When $S_{i}$ is empty, $c_{i}^{\textrm{loc}}$ is set to 0.0.
@@
The class certainty, $c_{i}^{\textrm{cls}}$ of reference detection $i$, is calculated using the entropy of the class probabilities $P(K)$:
@@
where $K$ is the set of classes. The entropy values are normalized to be in $[0,1]$ using the softmax function:
$\sigma_{i}(k)=\frac{e^{H_{i}(k)}}{\sum\limits_{k_{z}\in K}e^{H_{i}(k_{z})}},$ (11)
@@
The combined certainty, $c_{i}^{\textrm{comb}}$ of reference detection $i$, is calculated by taking the product of the occurrence, location and class certainty:
@@ 2.3 Data collection
Image data was collected from a grass field of around 0.75 hectare on the experimental field station of Wageningen University and Research near the city of Wageningen, The Netherlands. To allow repeatable measurements throughout the season, we simulate a weed detection application by distributing two different types of artificial plants on the grass field (Figure 5). The size of these plants is approximately 10 by 10 cm (top-down) and the minimum distance between two plants was set to 1m. Images were taken at four different dates (13 February 2024, 23 April 2024, 18 July 2024 and 1 Augustus 2024) to improve generalization of the detection network. Due to the availability of some parts of the field at some dates, the field shape and area slightly differ between the datasets.
Since weeds are non-uniformly distributed in fields (Cardina et al., 1997) and the spatial distribution of weeds can have large influence on the efficiency of an adaptive path planner (van Essen et al., 2024), we investigate the effect of different weed distributions. At each date, two datasets were recorded by arranging the plants in two different distributions: a uniform distribution and a clustered distribution (Figure 6 shows the distributions over the field on the dataset of 13 February 2024).
Images were aquired using a Matrice 300 RTK UAV equipped with Zenmuse P1 camera (DJI, Shenzhen, China) with 35mm focal length. An overlap of 70% side and 80% forward was used with an image resolution of 8192x5460 pixels. Images were recorded at 12m, 24m and 32m altitude. Additionally, eight ground control points (GCP) were placed in the field to improve georeferencing during creation of the orthomosaics, further explained in section 2.5. For evaluation, the location of the weeds and the GCPs were measured using a Topcon HiPer SR RTK-GNNS receiver (Topcon Positioning Systems, Inc., Tokyo, Japan) with around 1.0 - 1.5 cm accuracy.
@@
Figure 6: The measured location using RTK of the artificial plants distributed in an uniform (a) and clustered (b) distribution on the dataset of 13 February 2024. The colors indicate the different weed types.
@@ 2.4 Training of the detection network
For each date and distribution, 10% of the plants (6 plants per date and distribution) were randomly assigned to the training set and 5% to the validation set (3 plants per date and distribution). This leaves the majority of the plants unseen by the detection network, which was required for a fair evaluation of the adaptive path planner. Because each plant was visible in multiple images due to the overlap between the images, all images containing the training and validation plants were selected from the original UAV images at 12m, 24m and 32m altitude and manually annotated. To prevent data leakage between the training and validation dataset, in the training dataset, all plants that were not part of the training set were masked out by replacing the pixel values within the bounding box by a green color. Likewise, in the validation dataset, we masked out all plants that did not belong to the validation set. The detection network was trained on all datasets together, combining all dates, altitudes and distributions for generalization. The combined training dataset contained 1618 images with 2000 annotations and the combined validation dataset contained 889 images with 977 annotations respectively.
@@ 2.5 Offline simulation environment
To be able to do a large number of experiments and parameter optimization with the adaptive planner, we used the captured images to create georeferenced orthomosaics in order to get a realistic offline simulation environment. For each date and plant distribution, an orthomosaic was made from the 12m altitude images using Agisoft Metashape v2.1.0 (Agisoft LLC, St. Petersburg, Russia). The eight ground control points were used to improve georeferencing. In total, eight different orthomisaics were made using the high-quality settings.
For each waypoint, we calculated the field-of-view of the drone, rotated the orthomosaic using the UAV’s heading and cropped the field-of-view from the orthomosaic around the GNNS location of the UAV. The resulting image was resized to the original image size using linear interpolation. The generated images looked visually similar to the original images taken by the UAV.
@@ 2.6 Baseline
We compare the adaptive path planner with a coverage planner planned using Fields2Cover (Mier et al., 2023), covering the complete field at fixed altitude. Similarly as for the adaptive path planner, the row width and distance between two waypoints was set to 90% of the field-of-view.
@@ 2.7 Performance metrics
Both path planners were evaluated using the detection performance and flight path length. The locations of mapped objects with a certainty larger than $c_{\textrm{eval}}=0.5$ were compared to the ground truth using their GNNS location measured in the field. A mapped object was counted as a true positive (TP) when the class was equal to the ground truth class and the Euclidean distance with the ground truth location was smaller than 0.35m. When a mapped object had no ground truth location, it was counted as a false positive (FP), when an object was not present in the map, it was counted as a false negative (FN). The detection performance was quantified using the F1-score.
The path length was evaluated by summing the Euclidian distance between the waypoints. Since the area and shape of the datasets were not equal, we calculated the relative distance, $r_{k}^{\textrm{diff}}$, between the adaptive flight and the baseline coverage planner:
@@
where $l_{k}$ is the flight path length for the adaptive path planner on dataset $k$ and $l_{k}^{\textrm{cov}}$ is the length of the corresponding baseline coverage path at 12m altitude. When $r_{k}^{\textrm{diff}}$ is larger than 1, the adaptive flight path is longer than the baseline and when $r_{k}^{\textrm{diff}}$ is smaller than 1, its is shorter than the baseline.
@@ 2.8.1 Experiment 1: Relationship between detection certainty and errors
The certainty measure of a detection is an important part of the adaptive path planner, as it determines whether a low-altitude inspection is needed. Therefore, in order to make meaningful decisions, the certainty measure needs to provide separation between true and false positives. To investigate the usability of the different certainty measures (defined in 2.2.2) at different altitudes, we evaluated them on the captured images of 12m, 24m and 32m altitude for the 8 datasets. Errors made within 0.35m of the image border were ignored because the ground truth object could be outside the image due to inaccuracies in both the image GNNS location and measured ground truth object location which would result in an incorrect evaluation. For each certainty measure, we show the certainties for the TP and FP detections for the three altitudes. A large separability between the two indicates that the certainty measure can be used to differentiate between certain and uncertain detections. To quantify the separability between TP and FP detections using the certainty measure, we calculate the t-value for the alternative hypothesis that the mean confidence of the TPs, $\mu_{\textrm{tp}}$, is higher than the mean confidence of the false positive, $\mu_{\textrm{fp}}$, using Welch’s t-test (Welch, 1947):
@@
where $N$ is the sample size and $\sigma$ the standard deviation. The higher the value of $t$, the better the certainty measure can be used to differentiate between true and false positives. The certainty measure with the largest t-value for each altitude was used in experiment 2, 3 and 4.
@@ 2.8.2 Experiment 2: Parameter optimization
The altitude of the coverage flight path ($h_{\textrm{cov}}$), the acceptance confidence threshold ($c_{\textrm{accept}}$), and the rejection confidence threshold ($c_{\textrm{reject}}$) are important parameters for the adaptive path planner. These parameters are dependent on each other, for example, a higher coverage altitude works best with a low rejection confidence, because the detection certainty is lower at a higher altitude and a high rejection confidence will therefore reject most detections resulting in many false negatives. To study the influence of these parameters on the performance of the adaptive path planner, we ranged $h_{\textrm{cov}}$ from 12m to 48m in steps of 12m, $c_{\textrm{accept}}$ from 1.0 to 0.4 in steps of 0.2 and $c_{\textrm{reject}}$ with the values of 0.05, 0.2 and 0.4. All combinations are examined, except the combination of $c_{\textrm{accept}}=0.4$ and $c_{\textrm{reject}}=0.4$ as this combination would never result in the decision $d_{i}$ to inspect a detection (see section 2.1.5). We evaluated the F1-score and the relative difference in path length with the coverage flight path for all possible combinations of these parameters using the path planner on the clustered and uniform orthomosaics. The set of parameters that had both a good detection performance and a short path length were used in experiment 3 and 4. For all experiments, the inspection altitude, $h_{\textrm{inspect}}$, was set to 12m to match the minimum altitude of the UAV used during data collection. For comparison, we also evaluated the F1-score and relative flight path length for the coverage path planner for the same coverage altitudes ($h_{\textrm{cov}}$).
@@ 2.8.3 Experiment 3: Localization errors
Errors in the localization of the drone position and orientation directly influence the georeferencing of the detected objects. Georeferencing errors can results from errors in the position and altitude estimates by the UAV’s GNNS, and errors in estimates of roll, pitch and heading by the UAV’s IMU. To investigate the effect of the localization errors on the F1-score, we experimented with five levels of localization error. Table 1 defines the maximum position, altitude, gimbal roll, gimbal pitch and UAV heading errors for the different levels. The errors are sampled from a uniform distribution. The DJI M300 UAV used for the data collection had a positional error of around 1.5cm horizontal and vertical in the experiment field (provided by manufacturer) and an error in roll, pitch and heading of around 0.5 degree (average georeferencing error during the orthomosaic creation), which matches level ’good’ in Table 1. The errors in position and orientation were implemented by altering the cropped location and heading in the offline simulation environment. Because a higher localization error causes mismatching during the mapping of objects resulting in the creation of new objects in the map instead of updating existing ones (section 2.1.4), we increased the maximum distance between a mapped object and a detection, $d_{\textrm{dist}}$, to 0.5m, 0.9m, 1.8m, and 4.7m for the uncertainty levels good, decent, poor and very poor respectively. Level perfect keeps the default value of 0.35m. These values were computed by calculating the offset when all maximum errors, as defined in Table 1, occur simultaneously. Note that these maximum distances were only increased for the mapping, during evaluation we kept the maximum distance between a ground truth object and a detected object at the default value of 0.35m for a fair assessment. For each level, we compared the F1-score for the adaptive path planner with a coverage flight path at 12m, 24m, 36m and 48m altitude for both the clustered and uniform orthomosaics.
Table 1: Localization uncertainty levels used in experiment 3 expressed as offset in drone position, altitude and heading, and camera gimbal roll and pitch. The values are the minimum and maximum of the offset sampled from uniform distribution.

### v2

@@ 2.2.1 Monte-Carlo dropout
Dropout is a regularization technique in neural networks that randomly drops connections between neurons, and is regularly applied to avoid over-fitting during training by forcing the network to learn different ’routes’ through the network to get to the same outcome (Srivastava et al., 2014). Gal and Ghahramani (2016) introduced Monte-Carlo dropout during inference and showed that it can be used to assess model uncertainty. Each connection where dropout is applied has the probability of $q$ to be removed, and this process is repeated for $N$ runs in the network. When the network produces the same output during all runs while randomly removing some connections, it indicates that the network is certain. However, when every run results in a different output, the model is uncertain how to handle the input.
Figure 4 shows the architecture of YOLOv8, indicating the backbone, the neck and head of the network, and the several connections that pass through information from earlier layers in the network. The locations where dropout is applied are indicated with crosses. The dropout is applied between the neck and the head of the network, which enables us to reuse the output of the backbone and neck during inference and save computational time.
Figure 4: Schematic (simplified) overview of the YOLOv8 architecture with the locations indicated where Monte-Carlo dropout is applied between the neck and head of the network. Each ‘conv’ block contains a 2D convolution layer, batch normalization, and a sigmoid activation function. A ‘C2F’ block is a Cross-Stage Partial block with two convolutional layers. ‘SPPF’ indicates a Spatial Pyramid Pooling block, and ‘concat’ is a concatenation layer. The ‘detect’ block includes two ‘conv’ blocks followed by a 2D convolution layer to create the list of detections. See Jocher et al. (2023) for more details.
@@
where $D_{j,i}$, is the detection of object $i$ in the MCD run $j$. $D_{j,i}$ contains the bounding box coordinates, class probabilities, and confidence score. To combine the MCD detections with a the reference detection $D_{i}^{\textrm{ref}}$ for each forward pass $j$, we calculate the intersection-over-union (IoU) with all MCD detections $D_{j,i}^{\textrm{mc}}$:
$\textrm{IoU}(D_{i}^{\textrm{ref}},D_{j,i}^{\textrm{mc}})=\frac{|D_{i}^{\textrm{ref}}\cap D_{j,i}^{\textrm{mc}}|}{|D_{i}^{\textrm{ref}}\cup D_{j,i}^{\textrm{mc}}|}.$ (6)
When the IoU is larger than 0.5, $D_{j,i}^{\textrm{mc}}$ is added to the detection set $S_{i}$. Note that combining the reference detection with the MCD detections is only based on their bounding box coordinates and is class-independent.
@@ 2.2.2 Certainty calculation
The average YOLO certainty, $c_{i}^{\textrm{yolo}}$ of reference detection $i$ is calculated by taking the mean of the YOLO confidence values corresponding to the detection set $S_{i}$:
@@
$c_{i}^{\textrm{loc}}=\frac{1}{\|S_{i}\|}\cdot\sum_{D_{j,i}^{\textrm{mc}}\in S_{i}}\textrm{IoU}(D_{i}^{\textrm{ref}},D_{j,i}^{\textrm{mc}}).$ (9)
The localization certainty is based on the work of Blok et al. (2022). When $S_{i}$ is empty, $c_{i}^{\textrm{loc}}$ is set to 0.0.
@@
The class certainty of reference detection $i$ is calculated using the entropy, $H_{i}$:
@@
where $P_{j,i}(k)$ is the class probability for class $k\in K$ belonging to detection $j$ in the detection set $S_{i}$. The entropy values are normalized to be in $[0,1]$ using the softmax function, $\sigma_{i}(k)$:
$\sigma_{i}(k)=\frac{e^{H_{i}(k)}}{\sum\limits_{k_{z}\in K}e^{H_{i}(k_{z})}}.$ (11)
Finally, we calculate the class certainty, $c_{i}^{\textrm{cls}}$, by taking the maximum normalized certainty across all classes:
@@
The combined certainty, $c_{i}^{\textrm{comb}}$ of reference detection $i$, is calculated by taking the product of the occurrence, location, and class certainty:
@@ 2.3 Data collection
Image data was collected from a grass field of around 0.75 hectares on the experimental field station of Wageningen University and Research near the city of Wageningen, The Netherlands. To allow repeatable measurements throughout the season, we simulate a weed detection application by distributing two different types of artificial plants on the grass field (Figure 5). The size of these plants is approximately 10 by 10 cm (top-down), and the minimum distance between two plants was set to 1m. Images were taken on four different dates (13 February 2024, 23 April 2024, 18 July 2024, and 1 August 2024) to improve the generalization of the detection network. Due to the availability of some parts of the field on some dates, the field shape and area slightly differ between the datasets.
Since weeds are non-uniformly distributed in fields (Cardina et al., 1997) and the spatial distribution of weeds can have a large influence on the efficiency of an adaptive path planner (van Essen et al., 2024), we investigate the effect of different weed distributions. At each date, two datasets were recorded by arranging the plants in two different distributions: a uniform distribution and a clustered distribution (Figure 6 shows the distributions over the field on the dataset of 13 February 2024).
Images were acquired using a Matrice 300 RTK UAV equipped with a Zenmuse P1 camera (DJI, Shenzhen, China) with 35mm focal length. An overlap of 70% side and 80% forward was used with an image resolution of 8192x5460 pixels. Images were recorded at 12m, 24m, and 32m altitude. Additionally, eight ground control points (GCPs) were placed in the field to improve georeferencing during creation of the orthomosaics, further explained in section 2.5. For evaluation, the location of the weeds and the GCPs were measured using a Topcon HiPer SR RTK-GNSS receiver (Topcon Positioning Systems, Inc., Tokyo, Japan) with around 1.0 - 1.5 cm accuracy.
@@
Figure 6: The measured location using RTK of the artificial plants distributed in a uniform (a) and clustered (b) distribution on the dataset of 13 February 2024. The colors indicate the different weed types.
@@ 2.4 Training of the detection network
For each date and distribution, 10% of the plants (6 plants per date and distribution) were randomly assigned to the training set and 5% to the validation set (3 plants per date and distribution). This leaves the majority of the plants unseen by the detection network, which was required for a fair evaluation of the adaptive path planner. Because each plant was visible in multiple images due to the overlap between the images, all images containing the training and validation plants were selected from the original UAV images at 12m, 24m, and 32m altitude and manually annotated. To prevent data leakage between the training and validation datasets, in the training dataset, all plants that were not part of the training set were masked out by replacing the pixel values within the bounding box with a green color. Likewise, in the validation dataset, we masked out all plants that did not belong to the validation set. The detection network was trained on all datasets together, combining all dates, altitudes, and distributions for generalization. The combined training dataset contained 1618 images with 2000 annotations, and the combined validation dataset contained 889 images with 977 annotations, respectively.
@@ 2.5 Offline simulation environment
To be able to do a large number of experiments and parameter optimization with the adaptive planner, we used the captured images to create georeferenced orthomosaics in order to get a realistic offline simulation environment. For each date and plant distribution, an orthomosaic was made from the 12m altitude images using Agisoft Metashape v2.1.0 (Agisoft LLC, St. Petersburg, Russia). The eight ground control points were used to improve georeferencing. In total, eight different orthomosaics were made using the high-quality settings.
For each waypoint, we calculated the field-of-view of the drone, rotated the orthomosaic using the UAV’s heading and cropped the field-of-view from the orthomosaic around the GNSS location of the UAV. The resulting image was resized to the original image size using linear interpolation. The generated images looked visually similar to the original images taken by the UAV.
@@ 2.6 Baseline
We compare the adaptive path planner with a coverage planner planned using Fields2Cover (Mier et al., 2023), covering the complete field at fixed altitude. Similarly, as for the adaptive path planner, the row width and distance between two waypoints were set to 90% of the field-of-view.
@@ 2.7 Performance metrics
Both path planners were evaluated using the detection performance and flight path length. The locations of mapped objects with a certainty larger than $c_{\textrm{eval}}=0.5$ were compared to the ground truth using their GNSS location measured in the field. A mapped object was counted as a true positive (TP) when the class was equal to the ground truth class and the Euclidean distance from the ground truth location was smaller than 0.35m. When a mapped object had no ground truth location, it was counted as a false positive (FP), when an object was not present in the map, it was counted as a false negative (FN). The detection performance was quantified using the F1-score.
The path length was evaluated by summing the Euclidean distance between the waypoints. Since the area and shape of the datasets were not equal, we calculated the relative distance, $r_{k}^{\textrm{diff}}$, between the adaptive flight and the baseline coverage planner:
@@
where $l_{k}$ is the flight path length for the adaptive path planner on dataset $k$ and $l_{k}^{\textrm{cov}}$ is the length of the corresponding baseline coverage path at 12m altitude. When $r_{k}^{\textrm{diff}}$ is larger than 1, the adaptive flight path is longer than the baseline, and when $r_{k}^{\textrm{diff}}$ is smaller than 1, it is shorter than the baseline.
@@ 2.8.1 Experiment 1: Relationship between detection certainty and errors
The certainty measure of a detection is an important part of the adaptive path planner, as it determines whether a low-altitude inspection is needed. Therefore, in order to make meaningful decisions, the certainty measure needs to provide separation between true and false positives. To investigate the usability of the different certainty measures (defined in 2.2.2) at different altitudes, we evaluated them on the captured images of 12m, 24m, and 32m altitude for the 8 datasets. Errors made within 0.35m of the image border were ignored because the ground truth object could be outside the image due to inaccuracies in both the image GNSS location and measured ground truth object location, which would result in an incorrect evaluation. For each certainty measure, we show the certainties for the TP and FP detections for the three altitudes. A large separability between the two indicates that the certainty measure can be used to differentiate between certain and uncertain detections. To quantify the separability between TP and FP detections using the certainty measure, we calculate the t-value for the alternative hypothesis that the mean confidence of the TPs, $\mu_{\textrm{tp}}$, is higher than the mean confidence of the false positive, $\mu_{\textrm{fp}}$, using Welch’s t-test (Welch, 1947):
@@
where $N$ is the sample size and $\sigma$ the standard deviation. The higher the value of $t$, the better the certainty measure can be used to differentiate between true and false positives. The certainty measure with the largest t-value for each altitude was used in experiments 2, 3, and 4.
@@ 2.8.2 Experiment 2: Parameter optimization
The altitude of the coverage flight path ($h_{\textrm{cov}}$), the acceptance confidence threshold ($c_{\textrm{accept}}$), and the rejection confidence threshold ($c_{\textrm{reject}}$) are important parameters for the adaptive path planner. These parameters are dependent on each other; for example, a higher coverage altitude works best with a low rejection confidence, because the detection certainty is lower at a higher altitude, and a high rejection confidence will therefore reject most detections, resulting in many false negatives. To study the influence of these parameters on the performance of the adaptive path planner, we ranged $h_{\textrm{cov}}$ from 12m to 48m in steps of 12m, $c_{\textrm{accept}}$ from 1.0 to 0.4 in steps of 0.2 and $c_{\textrm{reject}}$ with the values of 0.05, 0.2 and 0.4. All combinations are examined, except the combination of $c_{\textrm{accept}}=0.4$ and $c_{\textrm{reject}}=0.4$ as this combination would never result in the decision $d_{i}$ to inspect a detection (see section 2.1.5). We evaluated the F1-score and the relative difference in path length with the coverage flight path for all possible combinations of these parameters using the path planner on the clustered and uniform orthomosaics. The set of parameters that had both a good detection performance and a short path length was used in experiments 3 and 4. For all experiments, the inspection altitude, $h_{\textrm{inspect}}$, was set to 12m to match the minimum altitude of the UAV used during data collection. For comparison, we also evaluated the F1-score and relative flight path length for the coverage path planner for the same coverage altitudes ($h_{\textrm{cov}}$).
@@ 2.8.3 Experiment 3: Localization errors
Errors in the localization of the drone position and orientation directly influence the georeferencing of the detected objects. Georeferencing errors can result from errors in the position and altitude estimates by the UAV’s GNSS, and errors in estimates of roll, pitch, and heading by the UAV’s IMU. To investigate the effect of the localization errors on the F1-score, we experimented with five levels of localization error. Table 1 defines the maximum position, altitude, gimbal roll, gimbal pitch, and UAV heading errors for the different levels. The errors are sampled from a uniform distribution. The DJI M300 UAV used for the data collection had a positional error of around 1.5cm horizontal and vertical in the experiment field (provided by manufacturer) and an error in roll, pitch and heading of around 0.5 degree (average georeferencing error during the orthomosaic creation), which matches level ’good’ in Table 1. The errors in position and orientation were implemented by altering the cropped location and heading in the offline simulation environment. Because a higher localization error causes mismatching during the mapping of objects resulting in the creation of new objects in the map instead of updating existing ones (section 2.1.4), we increased the maximum distance between a mapped object and a detection, $d_{\textrm{dist}}$, to 0.5m, 0.9m, 1.8m, and 4.7m for the uncertainty levels good, decent, poor and very poor respectively. Level perfect keeps the default value of 0.35m. These values were computed by calculating the offset when all maximum errors, as defined in Table 1, occur simultaneously. Note that these maximum distances were only increased for the mapping; during evaluation, we kept the maximum distance between a ground truth object and a detected object at the default value of 0.35m for a fair assessment. For each level, we compared the F1-score for the adaptive path planner with a coverage flight path at 12m, 24m, 36m and 48m altitude for both the clustered and uniform orthomosaics.
Table 1: Localization uncertainty levels used in experiment 3 expressed as offset in drone position, altitude, and heading, and camera gimbal roll and pitch. The values are the minimum and maximum of the offset sampled from a uniform distribution.